OpenStudy (anonymous):

Let \(r\) and \(s\) be two real numbers such that \[ r+s =-1 \\ r s =-1\] and let \(k \ge 1\) be an integer. Show that \[ r^k + s^k \] is a positive or negative integer.

OpenStudy (anonymous):

\((r+s)^k=\sum_{j=0}^{k }\binom{k}{j} r^js^{k-j}=r^k+s^k+\sum_{j=1}^{k-1 }\binom{k}{j} r^js^{k-j} \\ r^k+s^k=(r+s)^k-\sum_{j=1}^{k-1 }\binom{k}{j} r^js^{k-j} \\ \)

OpenStudy (anonymous):

\( r^k+s^k=(-1)^k-\sum_{j=1}^{k-1 }\binom{k}{j} r^j \frac{s^k}{s^j}\\= (-1)^k-\sum_{j=1}^{k-1 }\binom{k}{j} \left (\frac{r}{s} \right )^j s^k \\ \)

OpenStudy (anonymous):

rs=-1 r=-1/s

OpenStudy (anonymous):

r/s=-1/s^2

OpenStudy (solomonzelman):

`\(\large\color{black}{ \displaystyle \sum_{ n=1 }^{ \infty } ~ f(n)}\)` use this for inside the text sigma.

OpenStudy (anonymous):

\(\large r^k+s^k= (-1)^k- \sum_{j=1}^{k-1 }\binom{k}{j} (-1)^j s^{k-2j} \)

OpenStudy (anonymous):

hmm this is what i got so far xD

OpenStudy (anonymous):

when k is even 1- ( negative + positive + negative + positive +.....+negative ) if s>1 , second term also >1 1-negative is positive (case one ) if s<1 not sure (case 2) when k is odd -1 -(positive +negative+.....+ positive ) when s>1 then all term is negative (case 3) case 4 :- when s< 1 hmm

OpenStudy (anonymous):

hehe case 2 and 4 idk what to do in them :(

OpenStudy (anonymous):

Kinda wondered about induction when I saw this.

OpenStudy (anonymous):

induction for signs :|

OpenStudy (loser66):

I don't think induction for signs work just make few steps, we can see that it doesn't work if k odd, the basic step =-1, assume it works on k step, next is stuck: (r+k)^(k+1) = 1, positive.:(

OpenStudy (anonymous):

haha :D

OpenStudy (loser66):

Is there any outlet if we factor the term inside the sum as \(r^js^j (s^{k-2j})\) and then argue of j is odd or even?

OpenStudy (kainui):

When k is even, it's positive but that's all I know lol.

OpenStudy (loser66):

*if not of

OpenStudy (anonymous):

lol thats what i got , but wat if s<1 kai ?

OpenStudy (kainui):

Negative numbers squared are always positive =P

OpenStudy (anonymous):

lol imean <1 like 1/2 :P

OpenStudy (kainui):

Oh I see. We are showing that any two real numbers is not only positive or negative but that it is also an integer?

OpenStudy (anonymous):

its not necessary to be integer , but unlike integers real which <1 and >0 increase with decreasing power hmm

OpenStudy (anonymous):

so we wont know which term is max, specially with binomial coefficient or i might dint see it yet

OpenStudy (anonymous):

\[ (r+s)(r^k+s^k) = r^{k+1}+s^{k+1} +rs^{k}+sr^{k} = r^{k+1}+s^{k+1} -(r^{k-1}+s^{k-1})\\ \implies r^{k+1} +s^{k+1} = (r+s)(r^k+s^k)+(r^{k-1}+s^{k-1}) \]We assume that by indictive hypothesis that \(p(k-1)\land p(k)\).\[ r^{k+1} +s^{k+1} = -m+n \in\mathbb Z \]Then we have to show the base case?

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

inductive hypothesis

OpenStudy (anonymous):

cool !

OpenStudy (anonymous):

I usually am not this clever... I just guessed.

OpenStudy (anonymous):

wait the trick is by being integer xD

OpenStudy (anonymous):

wait lol all above is not needed :|

OpenStudy (anonymous):

s+r=-1 and rs=-1 is solvable equation

OpenStudy (anonymous):

we only need to find set on solution in R :(

OpenStudy (anonymous):

so from wolfram http://www.wolframalpha.com/input/?i=r%2Bs%3D-1+%2C+rs%3D-1

OpenStudy (anonymous):

and remember this post http://openstudy.com/users/marki#/updates/54987c7ae4b0e4319f8bf7ae

OpenStudy (anonymous):

what i was thinking before :|

OpenStudy (anonymous):

Here is an easy proof: r and s are solutions of \[ x^2 + x-1=0\] hence \[ r^2 = 1 -r\\ r^k = r^{k-2} - r^{k-1}\\ s^2 = 1 -r\\ s^k = s^{k-2} - s^{k-1}\\ r^k+ s^k = r^{k-2}+ s^{k-2} - (r^{k-1}+s^{k-1})\\ \] And use induction. Noticing that \[ r+s=-1\\ r^2+s^2=(r+s)^2 - 2 r s=1+2=3 \]

OpenStudy (anonymous):

@marki @ganeshie8

OpenStudy (anonymous):

This problem was inspired by http://openstudy.com/users/marki#/updates/54987c7ae4b0e4319f8bf7ae where it was asked that \[ r^6+ s^6 =18 \] This is a kind of generalization of it. @campbell_st

OpenStudy (anonymous):

\[ s^2 = 1- s\] in my solution above.

OpenStudy (anonymous):

nice !

ganeshie8 (ganeshie8):

Clever xD I think this problem is equivalent to proving \((-1+\sqrt{5})^n + (-1-\sqrt{5})^n\) is divisible by \(2^n\) for all integers \(n\ge 1\)

OpenStudy (anonymous):

Yes it is @ganeshie8